OpenStudy (anonymous):

ok here we go again 2/3 t -5=1 what is the solution t=?

jimthompson5910 (jim_thompson5910):

You can multiply both sides by 3 to clear out the fraction

jimthompson5910 (jim_thompson5910):

(2/3)t - 5 = 1 3*[(2/3)t - 5] = 3*1 ... Multiply both sides by 3. 3*(2/3)t - 3*5 = 3 2t - 15 = 3

jimthompson5910 (jim_thompson5910):

so (2/3)t - 5 = 1 turns into 2t - 15 = 3

OpenStudy (anonymous):

okay you make it look so easy wow

jimthompson5910 (jim_thompson5910):

I'm glad it's clicking better now

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

so again the rule is: if you have fractions in an equation, multiply both sides by the LCD (lowest common denominator) to clear them out

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

yes thank you so much yoi where a great help

jimthompson5910 (jim_thompson5910):

I'm glad I helped out

OpenStudy (anonymous):

@_@ jim_thompson5910